# No-Rank Tensor Decomposition Using Metric Learning

### v1

@@ Abstract.
Tensor decomposition faces fundamental challenges in analyzing high-dimensional data, where traditional methods based on reconstruction and fixed-rank constraints often fail to capture semantically meaningful structures. This paper introduces a no-rank tensor decomposition framework grounded in metric learning, which replaces reconstruction objectives with a discriminative, similarity-based optimization. The proposed approach learns data-driven embeddings by optimizing a triplet loss with diversity and uniformity regularization, creating a feature space where distance directly reflects semantic similarity. We provide theoretical guarantees for the framework’s convergence and establish bounds on its metric properties. Evaluations across diverse domains—including face recognition (LFW, Olivetti), brain connectivity analysis (ABIDE), and simulated data (galaxy morphology, crystal structures)—demonstrate that our method outperforms baseline techniques, including PCA, t-SNE, UMAP, and tensor decomposition baselines (CP and Tucker). Results show substantial improvements in clustering metrics (Silhouette Score, Davies-Bouldin Index, Calinski-Harabasz Index, Separation Ratio, Adjusted Rand Index, Normalized Mutual Information) and reveal a fundamental trade-off: while metric learning optimizes global class separation, it deliberately transforms local geometry to align with semantic relationships. Crucially, our approach achieves superior performance with smaller training datasets compared to transformer-based methods, offering an efficient alternative for domains with limited labeled data. This work establishes metric learning as a paradigm for tensor-based analysis, prioritizing semantic relevance over pixel-level fidelity while providing computational advantages in data-scarce scenarios.
@@ 1. Introduction
Tensor decomposition and representation learning are two fundamental paradigms for extracting meaningful structure from multi-dimensional data. Traditional tensor decomposition methods, such as CANDECOMP/PARAFAC (CP) [kolda2009tensor] and Tucker decomposition [tucker1966some], provide mathematically elegant approaches for factorizing tensors into interpretable components.
@@
While these methods offer strong uniqueness guarantees [kruskal1977three] and interpretability, they require pre-specification of rank parameters $(R,R_{1},\dots,R_{N})$ and are inherently limited to multi-linear relationships.
Representation learning [bengio2013representation] provides a complementary approach, learning useful data representations through neural networks without explicit structural constraints. Deep autoencoders, for instance, learn an encoding function $f_{\theta}$ mapping input data to latent codes $\mathbf{z}$ and a decoding function $g_{\phi}$ reconstructing the original input:
@@
Recent work has begun bridging these domains, with neural networks enhancing tensor decompositions [liu2022tensor] and implicit neural representations demonstrating remarkable compression capabilities [tancik2020fourier]. However, the fundamental challenge of rank selection persists across traditional methods.
@@ 2. Related Work
Tensor decomposition methods have long been the workhorse for analyzing multi-way data. The Canonical Polyadic (CP) decomposition [hitchcock1927expression, kolda2009tensor] and the Tucker decomposition [tucker1966some] are the two most foundational approaches. Both methods impose a strict low-rank constraint on the data, seeking an approximation $\hat{\mathcal{X}}$ that minimizes the reconstruction error $\mathcal{X}-\hat{\mathcal{X}}||_{F}^{2}$. While effective for data compression and denoising, this reconstructive objective is not inherently aligned with discriminative tasks like classification or clustering. The requirement to pre-define the rank or multilinear rank is a significant limitation, as the intrinsic data complexity is often unknown and may not be well-represented by a low-rank model [acar2011scalable].
Subsequent advancements, such as tensor-train decompositions [oseledets2011tensor] and non-negative tensor factorizations [cichocki2009nonnegative], have improved scalability and interpretability but have largely remained within the reconstructive paradigm. These methods are fundamentally linear and struggle to capture the complex, non-linear manifolds on which high-dimensional data often resides.
@@
To address non-linearity, a separate lineage of work focused on manifold learning and non-linear dimensionality reduction [bagherian2023classical]. Techniques such as Isomap [tenenbaum2000global], Locally Linear Embedding (LLE) [roweis2000nonlinear], and Laplacian Eigenmaps [belkin2003laplacian] aim to preserve geometric properties of the data. More recently, t-Distributed Stochastic Neighbor Embedding (t-SNE) [maaten2008visualizing] and Uniform Manifold Approximation and Projection (UMAP) [mcinnes2018umap] have become standards for visualization, excelling at preserving local neighborhood structures.
@@
The advent of deep learning catalyzed a shift towards learning representations directly optimized for a specific task. A pivotal development in this space is metric learning [kulis2013metric, bellet2015survey], which aims to learn a distance function that reflects semantic similarity. The contrastive loss [hadsell2006dimensionality] and, more influentially, the triplet loss [schroff2015facenet] provided a powerful framework for this. By pulling an anchor sample closer to a positive sample than to a negative sample by a margin, these losses directly optimize the embedding space for discrimination.
This paradigm has driven state-of-the-art performance in face recognition [schroff2015facenet, wang2018cosface], image retrieval [faghri2018vse], and person re-identification [hermans2017defense]. This work draws direct inspiration from these successes but adapts the triplet loss framework to the problem of tensor decomposition, moving beyond its typical application in computer vision.
To prevent pathological solutions like dimensional collapse, recent work has emphasized the importance of regularization. Wang and Isola [wang2020understanding] identified alignment and uniformity as key properties of effective representations, which has led to the use of uniformity losses [wang2020understanding] and diversity penalties on the embedding correlation matrix [bansal2020canonical]. The proposed framework incorporates these insights to ensure a well-structured and effective embedding space.
@@
There is a growing recognition of the limitations of traditional methods for data analysis. In domains like astronomy, methods for galaxy classification have evolved from manual taxonomy [hubble1926extragalactic] to machine learning approaches using hand-crafted features [dieleman2015rotation] and, more recently, deep convolutional networks [walmsley2022galaxy]. Similarly, in materials science, the analysis of crystal structures has been tackled with symmetry-based descriptors [isayev2017universality] and graph neural networks [xie2021crystal].
However, the application of deep metric learning in these contexts is still nascent. While some studies have used contrastive learning for spectroscopic data [stein2022spectroscopic] or medical imaging [sun2022metric], a generalized framework that replaces the core principles of tensor decomposition for tensor data is lacking. Existing approaches often treat metric learning as a separate module rather than as a fundamental alternative to decomposition.
Built on the previous works [bagherian2021coupled, bagherian2022bilevel, bagherian2024tensor], the proposed no-rank tensor decomposition framework synthesizes ideas from these disparate fields. We reframe the problem of tensor analysis from one of reconstruction to one of discrimination, drawing on the power of triplet-based metric learning. Unlike traditional tensor methods, we impose no explicit rank constraints and leverage deep non-linear networks to capture complex data manifolds. Unlike unsupervised manifold learning, we directly optimize for semantic similarity using label information. And unlike standard metric learning applications, the proposed method is positioned as a direct, end-to-end replacement for tensor decomposition in the data pipeline, a novel contribution with significant practical implications.
@@ 3.1. Theoretical Framework
Traditional tensor decomposition methods, such as Canonical Polyadic (CP) and Tucker decompositions, impose strict rank constraints that may not align with the intrinsic geometry of high-dimensional data. We propose a no-rank tensor decomposition framework based on metric learning, which learns data-driven similarity structures without explicit rank constraints.
@@
The model is trained to pull the anchor and positive together while pushing the anchor and negative apart, leading to an embedding space where semantic similarity is inversly proportional to the distance in embedding space. This triplet formulation provides a flexible framework for Semantic Structure Learning, directly optimizing for similarity relationships rather than reconstruction error.

### v2

# arXiv:2511.01816v2[cs.LG] 27 Feb 2026

@@ Abstract.
Tensor decomposition of high-dimensional data often struggles to capture semantically or physically meaningful structures, particularly when relying on reconstruction objectives and fixed-rank constraints. We introduce a no-rank tensor decomposition framework based on metric learning, which replaces reconstruction objectives with a similarity-driven optimization. By combining a triplet loss with diversity and uniformity regularization, the method learns embeddings where distances naturally reflect semantic and physical relationships, supported by theoretical guarantees on convergence and metric properties.
We evaluate the approach on diverse datasets, including face recognition (LFW, Olivetti), brain connectivity (ABIDE), and simulated physical systems (galaxies, crystals). In comprehensive comparisons against classical methods (PCA, t-SNE, UMAP), tensor decompositions (CP, Tucker, t-SVD), and deep learning models (VAE, DEC, transformer-based embeddings), our method produces embeddings that preserve physically and semantically relevant relationships and achieve competitive clustering performance. While transformers often excel in predictive accuracy on large datasets, our method provides interpretable embeddings and remains effective in small-data regimes where transformer training may be infeasible.
This work establishes metric learning as a principled paradigm for tensor analysis, emphasizing physical interpretability and semantic relevance over pixel-level reconstruction, and offering an efficient and robust alternative in data-scarce scientific domains.
@@ 1. Introduction
Tensor decomposition and representation learning are two fundamental paradigms for extracting meaningful structure from multi-dimensional data. Traditional tensor decomposition methods, such as CANDECOMP/PARAFAC (CP) [26] and Tucker decomposition [45], provide mathematically elegant approaches for factorizing tensors into interpretable components.
@@
While these methods offer strong uniqueness guarantees [27] and interpretability, they require pre-specification of rank parameters $(R,R_{1},\dots,R_{N})$ and are inherently limited to multi-linear relationships.
Representation learning [7] provides a complementary approach, learning useful data representations through neural networks without explicit structural constraints. Deep autoencoders, for instance, learn an encoding function $f_{\theta}$ mapping input data to latent codes $\mathbf{z}$ and a decoding function $g_{\phi}$ reconstructing the original input:
@@
Recent work has begun bridging these domains, with neural networks enhancing tensor decompositions and implicit neural representations demonstrating remarkable compression capabilities [43]. However, the fundamental challenge of rank selection persists across traditional methods.
@@ 2. Related Work
Tensor decomposition methods have long been the workhorse for analyzing multi-way data. The Canonical Polyadic (CP) decomposition [20] and the Tucker decomposition [45] are the two most foundational approaches. Both methods impose a strict low-rank constraint on the data, seeking an approximation $\hat{\mathcal{X}}$ that minimizes the reconstruction error $\mathcal{X}-\hat{\mathcal{X}}||_{F}^{2}$. While effective for data compression and denoising, this reconstructive objective is not inherently aligned with discriminative tasks like classification or clustering. The requirement to pre-define the rank or multilinear rank is a significant limitation, as the intrinsic data complexity is often unknown and may not be well-represented by a low-rank model [1].
Subsequent advancements, such as tensor-train decompositions [36] and non-negative tensor factorizations [10], have improved scalability and interpretability but have largely remained within the reconstructive paradigm. These methods are fundamentally linear and struggle to capture the complex, non-linear manifolds on which high-dimensional data often resides.
Recent work has enhanced tensor decompositions by incorporating structured operators. For instance, authors in [58] proposed a tensor convolution-like low-rank dictionary for high-dimensional image representation, improving the capture of local spatial correlations. Nevertheless, such methods remain reconstructive and rely on explicit low-rank constraints, limiting their suitability for discriminative tasks.
Other modern tensor decomposition approaches include t-SVD-based frameworks, which generalize matrix SVD to third-order tensors using the t-product. These methods have been applied to representation learning via tensor neural networks under transformed low-rank constraints, [49], and to manifold learning using t-product geometry [50]. Similar to CP, Tucker, and convolutional low-rank dictionaries, t-SVD approaches rely on explicit low-rank structures and linear transformations, and are primarily reconstructive. In contrast, our no-rank metric learning framework learns embeddings without imposing rank constraints, directly optimizing semantic similarity in a non-linear, discriminative manner.
@@
To address non-linearity, a separate lineage of work focused on manifold learning and non-linear dimensionality reduction [2]. Techniques such as Isomap [44], Locally Linear Embedding (LLE) [39], and Laplacian Eigenmaps [6] aim to preserve geometric properties of the data. More recently, t-Distributed Stochastic Neighbor Embedding (t-SNE) [46] and Uniform Manifold Approximation and Projection (UMAP) [32] have become standards for visualization, excelling at preserving local neighborhood structures.
@@
The advent of deep learning catalyzed a shift towards learning representations directly optimized for a specific task. A pivotal development in this space is metric learning [28], which aims to learn a distance function that reflects semantic similarity. The contrastive loss [16] and, more influentially, the triplet loss [40] provided a powerful framework for this. By pulling an anchor sample closer to a positive sample than to a negative sample by a margin, these losses directly optimize the embedding space for discrimination.
This paradigm has driven state-of-the-art performance in face recognition [51], image retrieval [15], and person re-identification [18]. This work draws direct inspiration from these successes but adapts the triplet loss framework to the problem of tensor decomposition, moving beyond its typical application in computer vision.
To prevent pathological solutions like dimensional collapse, recent work has emphasized the importance of regularization. Wang and Isola [52] identified alignment and uniformity as key properties of effective representations, which has led to the use of uniformity losses and diversity penalties on the embedding correlation matrix [29]. The proposed framework incorporates these insights to ensure a well-structured and effective embedding space.
@@
There is a growing recognition of the limitations of traditional methods for data analysis. In domains like astronomy, methods for galaxy classification have evolved from manual taxonomy [23] to machine learning approaches using hand-crafted features [13] and, more recently, deep convolutional networks [48]. Similarly, in materials science, the analysis of crystal structures has been tackled with symmetry-based descriptors [24] and graph neural networks [57].
However, the application of deep metric learning in tensor-structured scientific data remains nascent. While contrastive and metric learning techniques have been successfully applied in domains such as medical imaging [9] and scientific signal representation learning [35], a generalized framework that replaces the core principles of tensor decomposition for multiway data is still lacking. Existing approaches typically treat metric learning as an auxiliary representation module rather than as a fundamental alternative to tensor decomposition.
Built on the previous works [3, 4, 5], the proposed no-rank tensor decomposition framework synthesizes ideas from these disparate fields. We reframe the problem of tensor analysis from one of reconstruction to one of discrimination, drawing on the power of triplet-based metric learning. Unlike traditional tensor methods, we impose no explicit rank constraints and leverage deep non-linear networks to capture complex data manifolds. Unlike unsupervised manifold learning, we directly optimize for semantic similarity using label information. And unlike standard metric learning applications, the proposed method is positioned as a direct, end-to-end replacement for tensor decomposition in the data pipeline, a contribution with practical implications.
@@ 3.1. Theoretical Framework
Traditional tensor decomposition methods, such as Canonical Polyadic (CP) and Tucker decompositions, impose strict rank constraints that may not align with the intrinsic geometry of high-dimensional data. Tensor rank is NP-hard and unstable to compute [19]. Therefore, we propose a no-rank tensor decomposition framework based on metric learning, which learns data-driven similarity structures without explicit rank constraints.
Given a learned embedding matrix $\mathbf{Z}\in\mathbb{R}^{n\times d}$ where rows correspond to data samples and columns to embedding dimensions, the effective rank quantifies the number of significant, linearly independent dimensions in the learned representation. Let $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{d}\geq 0$ be the singular values of $\mathbf{Z}$. The following standard measures of effective rank are equivalent:
  1. (i)
    Explained variance threshold: For $\epsilon\in[0,1)$,
    $\text{erank}_{\epsilon}(\mathbf{Z})=\min\left\{k:\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sum_{i=1}^{d}\sigma_{i}^{2}}\geq 1-\epsilon\right\}.$
  2. (ii)
    Spectral decay ratio: For $\tau\in(0,1]$,
    $\text{erank}_{\tau}(\mathbf{Z})=\max\left\{k:\sigma_{k}\geq\tau\cdot\sigma_{1}\right\}.$
  3. (iii)
    Entropy-based effective rank:
    $\text{erank}_{\text{ent}}(\mathbf{Z})=\exp\left(-\sum_{i=1}^{d}p_{i}\log p_{i}\right),\quad\text{where }p_{i}=\frac{\sigma_{i}^{2}}{\sum_{j=1}^{d}\sigma_{j}^{2}}.$
In the context of no-rank tensor decomposition, these measures capture the implicit dimensionality learned through metric optimization. When embedding dimensions become decorrelated (a property enforced by our diversity regularization), the columns of $\mathbf{Z}$ become approximately orthonormal and the nonzero singular values concentrate around one. Consequently, all three effective rank measures converge to the embedding dimension $d$.
Definition 1 (No-Rank Tensor Decomposition).
Given an $N$-th order tensor $\mathcal{X}\in\mathbb{R}^{I_{1}\times\cdots\times I_{N}}$, a no-rank tensor decomposition learns embedding functions $f^{(n)}:\mathbb{R}^{I_{\setminus n}}\to\mathbb{R}^{d}$ (possibly shared across modes) that map mode-$n$ fibers to semantic embeddings.
For each mode $n=1,\dots,N$, let $\mathbf{x}_{i_{n}}^{(n)}$ be the mode-$n$ fiber at index $i_{n}$, and define $\mathbf{z}_{i_{n}}^{(n)}=f^{(n)}(\mathbf{x}_{i_{n}}^{(n)})\in\mathbb{R}^{d}$.
The decomposition is characterized by (i) an implicit similarity tensor $\mathcal{S}\in\mathbb{R}^{I_{1}\times\cdots\times I_{N}}$ with entries
$\mathcal{S}_{i_{1},\dots,i_{N}}=\langle\mathbf{z}_{i_{1}}^{(1)},\mathbf{z}_{i_{2}}^{(2)},\dots,\mathbf{z}_{i_{N}}^{(N)}\rangle.$
(ii) the effective rank of the decomposition is defined by the effective rank of the embedding matrices $\mathbf{Z}^{(n)}=[\mathbf{z}_{1}^{(n)^{\top}},\dots,\mathbf{z}_{I_{n}}^{(n)^{\top}}]^{\top}$, determined implicitly through optimization rather than prescribed a priori.
This notion of decomposition is formalized in Proposition 5, which shows that the induced similarity tensor admits a CP-style factorization whose effective rank equals the embedding dimension.
@@
The model is trained to pull the anchor and positive together while pushing the anchor and negative apart, leading to an embedding space in which semantic similarity is inversely related to Euclidean distance. This triplet formulation provides a flexible framework for Semantic Structure Learning, directly optimizing for similarity relationships rather than reconstruction error.
3.2. Connection to Tensor Rank
While avoiding explicit rank constraints, our method maintains a connection to tensor algebra through the embedding dimension. We now formalize the role of the diversity regularization term in determining the effective dimensionality of the learned representation.
Remark 2 (Effective vs. Algebraic Rank).
We emphasize that this notion of rank is effective rather than algebraic: it characterizes the number of active, linearly independent latent components induced by the embedding geometry and optimization dynamics, rather than the minimal CP rank in the classical sense.
Proposition 3 (Embedding Dimension as Effective Rank).
Let $\mathbf{Z}=[\mathbf{z}_{1}^{\top},\dots,\mathbf{z}_{n}^{\top}]^{\top}\in\mathbb{R}^{n\times d}$ be the learned embedding matrix, where rows correspond to data samples and columns to embedding dimensions. Let each column be $\ell_{2}$-normalized, i.e., $\|\mathbf{z}^{(k)}\|_{2}=1$ for all $k=1,\dots,d$. Define the correlation matrix $\mathbf{C}=\mathbf{Z}^{\top}\mathbf{Z}\in\mathbb{R}^{d\times d}$ and the diversity loss
$\mathcal{L}_{\text{div}}=\frac{1}{d(d-1)}\sum_{i\neq j}|\mathbf{C}_{ij}|.$ (4)
If $\mathcal{L}_{\text{div}}\to 0$ during optimization, then $\mathbf{C}\to\mathbf{I}_{d}$ and $\mathbf{Z}$ has full column rank. Consequently, the embedding dimension $d$ serves as an effective rank determined by optimization.
Proof.
By construction, the correlation matrix $\mathbf{C}=\mathbf{Z}^{\top}\mathbf{Z}$ is symmetric and positive semidefinite. Since each column of $\mathbf{Z}$ is $\ell_{2}$-normalized, the diagonal entries satisfy
$\mathbf{C}_{ii}=\langle\mathbf{z}^{(i)},\mathbf{z}^{(i)}\rangle=1\quad\text{for all }i.$
The diversity loss penalizes all off-diagonal correlations. Hence
$\mathcal{L}_{\text{div}}\to 0\quad\Rightarrow\quad\mathbf{C}_{ij}\to 0\text{ for all }i\neq j.$
Therefore,
$\mathbf{C}\to\mathbf{I}_{d}.$
Since $\mathbf{Z}^{\top}\mathbf{Z}$ becomes the identity, the columns of $\mathbf{Z}$ become asymptotically orthonormal, and $\mathbf{Z}$ has full column rank:
$\operatorname{rank}(\mathbf{Z})=d.$
Remark 4.
As $\mathbf{C}\to\mathbf{I}_{d}$, the nonzero singular values of $\mathbf{Z}$ concentrate around one. Consequently, the effective rank measures satisfy
$\text{erank}_{\epsilon}(\mathbf{Z})\to d,\quad\text{erank}_{\tau}(\mathbf{Z})\to d,\quad\text{erank}_{\text{ent}}(\mathbf{Z})\to d,$
for any $\epsilon<1$ and $\tau\leq 1$. Thus, the optimization learns representations with maximal effective rank.
The triplet-based objective induces an implicit rank structure in the learned representations. By minimizing intra-class distances, the embeddings of samples within the same semantic class concentrate on low-dimensional subspaces, while maximizing inter-class distances enforces separation across distinct subspaces. Through diversity and uniformity regularization, the overall embedding adapts its effective dimensionality to the intrinsic complexity of the data, without requiring explicit rank selection. This property is particularly advantageous in applications where semantic organization is more relevant than exact reconstruction, such as medical imaging, retrieval, and high-dimensional pattern analysis.
3.3. Metric-Induced Tensor Factorization
The previous proposition established how diversity regularization controls the effective rank of the embedding matrix. We now extend this analysis to the full tensor setting, showing that the induced similarity tensor $\mathcal{S}$ admits a classical CP decomposition whose rank is implicitly determined by the metric learning objective. This result formally connects our no-rank framework to traditional tensor algebra.
Proposition 5 (Metric Learning Induces CP Structure).
Let $\mathcal{X}\in\mathbb{R}^{I_{1}\times\cdots\times I_{N}}$ be an $N$-th order tensor, and let embeddings be learned through the metric learning objective with triplet loss $\mathcal{L}_{\text{triplet}}$ and diversity regularization $\mathcal{L}_{\text{div}}$ (Eq. (4)).
After training, the learned embedding function $f$ produces mode-$n$ embeddings:
$\mathbf{z}_{i_{n}}^{(n)}=f(\mathcal{X}_{:\cdots i_{n}\cdots:})\in\mathbb{R}^{d}.$
Define the similarity tensor $\mathcal{S}$ induced by the metric learning objective:
$\mathcal{S}_{i_{1},\dots,i_{N}}=\left\langle\mathbf{z}_{i_{1}}^{(1)},\mathbf{z}_{i_{2}}^{(2)},\dots,\mathbf{z}_{i_{N}}^{(N)}\right\rangle.$
Then:
  1. (a)
    $\mathcal{S}$ admits a CP decomposition with exactly $d$ components:
    $\mathcal{S}=\sum_{r=1}^{d}\mathbf{z}_{:r}^{(1)}\circ\mathbf{z}_{:r}^{(2)}\circ\cdots\circ\mathbf{z}_{:r}^{(N)},$
    where $\mathbf{z}_{:r}^{(n)}=[z^{(n)}_{1r},\dots,z^{(n)}_{I_{n}r}$$^{\top}$.
  2. (b)
    The factor matrices $\mathbf{Z}^{(n)}=[\mathbf{z}_{1}^{(n)},\dots,\mathbf{z}_{I_{n}}^{(n)}]^{\top}\in\mathbb{R}^{I_{n}\times d}$ are determined by the metric learning optimization:
    $\mathbf{Z}^{(n)}=\arg\min_{\mathbf{Z}}\left[\mathcal{L}_{\text{triplet}}(\mathbf{Z})+\lambda\mathcal{L}_{\text{div}}(\mathbf{Z})\right],$
    where $\lambda>0$ balances the objectives.
  3. (c)
    When $\mathcal{L}_{\text{div}}\to 0$, each $\mathbf{Z}^{(n)}$ has full column rank, making $d$ the effective CP rank of $\mathcal{S}$. This rank emerges implicitly from the interplay between triplet loss (shaping semantic geometry) and diversity regularization (controlling dimensionality).
Proof.
The CP structure follows algebraically from expanding the inner product:
$\langle\mathbf{z}_{i_{1}}^{(1)},\dots,\mathbf{z}_{i_{N}}^{(N)}\rangle=\sum_{r=1}^{d}z^{(1)}_{i_{1}r}\cdots z^{(N)}_{i_{N}r}$.
The key observation is that the factors $\mathbf{Z}^{(n)}$ are not free variables but outcomes of metric learning. Formally, let $\Theta$ parameterize $f$. The optimization:
$\min_{\Theta}\mathbb{E}_{(a,p,n)\sim\mathcal{T}}\left[\|\mathbf{z}_{a}-\mathbf{z}_{p}\|^{2}-\|\mathbf{z}_{a}-\mathbf{z}_{n}\|^{2}+\alpha\right]_{+}+\lambda\mathcal{L}_{\text{div}}(\mathbf{Z})$
determines $\mathbf{Z}^{(n)}=f(\mathcal{X}^{(n)};\Theta^{*})$, where $\mathcal{X}^{(n)}$ denotes mode-$n$ slices. Unlike classical CP where $\mathbf{Z}^{(n)}$ minimize $\|\mathcal{X}-\sum_{r}\mathbf{z}_{:r}^{(1)}\circ\cdots\circ\mathbf{z}_{:r}^{(N)}\|$, here they minimize semantic distance violations.
When $\mathcal{L}_{\text{div}}\to 0$, we have $(\mathbf{Z}^{(n)})^{\top}\mathbf{Z}^{(n)}\to\mathbf{I}_{d}$, implying $\text{rank}(\mathbf{Z}^{(n)})=d$. Thus the CP representation uses all $d$ components non-degenerately, making $d$ the effective rank. ∎
Remark 6 (Interpretation of Metric-Induced Tensor Factorization).
This result formalizes the connection between metric learning and tensor decomposition. While classical CP and Tucker methods decompose tensor entries by minimizing reconstruction error, our framework decomposes the similarity tensor $\mathcal{S}$ induced by embeddings optimized for semantic relationships.
The key distinction is that factors $\mathbf{Z}^{(n)}$ are not obtained by solving $\min\|\mathcal{X}-\sum_{r}\mathbf{z}_{:r}^{(1)}\circ\cdots\circ\mathbf{z}_{:r}^{(N)}\|$, but rather emerge from metric learning: the triplet loss shapes embedding geometry so inner products encode semantic similarity, while diversity regularization controls dimensionality. Consequently, the CP structure is not imposed a priori but emerges from the optimization process, with effective rank determined implicitly rather than chosen explicitly.